solve this physics question on electric field -the center-to-center separation of two tiny, spherical water drops, with identical charges of -1 × 10^-16C is 1cm, calculate the magnitude of the electrostatic force acting between them and the excess electrons on each drop, giving it its charge imbalance?
1
Expert's answer
2019-09-16T10:20:47-0400

Question (a)


To calculate the electrostatic force, Coulomb's Law is applied.


"F_{E}=K\frac{q_{1}q_{2}}{r^{2}}"


Where:

  • The coulomb constant is "k=9*10^{9}Nw\frac{m^{2}}{C^{2}}"
  • The magnitude of electric charges "q_{1}=q_{2}=-1*10^{-16}C"
  • The distance between charges is "r=1cm*\frac{1m}{100cm}=0.01m"



Numerically evaluating


"F_{E}=9*10^{9}Nw\frac{m^{2}}{C^{2}}*\frac{-1*10^{-16}C*-1*10^{-16}C}{(0.01m)^{2}}\\ F_{E}=9*10^{-19}Nw"


The electrostatic force between the two drop is


"\boxed{F_{E}=9*10^{-19}Nw}"


Question (b)


The charge of an electron is:"e^{-}=-1.60*10^{-19}C"


The net charge in each drop is "q=-1*10^{-16}C"


Therefore the number of electrons in each sphere is


"N_{e^{-}}=\frac{-1*10^{-16}C}{-1.60*10^{-19} C}=625electrons"


The number of electrons in each drop is.

"\boxed{N_{e^{-}}=625electrons}"